# Decision Quotient: A Regime-Sensitive Complexity Theory of Exact Relevance Certification

### 16 Mar 2026 v1

@@ Abstract
Which coordinates of a decision problem can be hidden without changing the decision, and what is the coarsest exact abstraction that preserves all decision-relevant distinctions? We study this as an exact relevance-certification problem organized around the optimizer quotient. We classify how hard it is to certify this structure across three settings: static (counterexample exclusion), stochastic (conditioning and expectation), and sequential (temporal structure). In the static regime, sufficiency collapses to relevance containment, so minimum sufficiency is coNP-complete. In the stochastic regime, preservation and decisiveness separate: preservation is polynomial-time under explicit-state encoding with bridge theorems to static sufficiency and the optimizer quotient, while decisiveness is PP-hard under succinct encoding with anchor and minimum variants in $\textsf{NP}^{\textsf{PP}}$. In the sequential regime, all queries are PSPACE-complete. We also prove an encoding-sensitive contrast between explicit-state tractability and succinct-encoding hardness, derive an integrity-competence trilemma, and identify tractable subcases across two tiers. A Lean 4 artifact mechanically verifies the optimizer-quotient universal property, main reductions, and finite decider core.
footnotetext: © 2026 Tristan Simas. This work is licensed under $C$ BY 4.0. License: https://creativecommons.org/licenses/by/4.0/
1 Introduction
Which coordinates of a decision problem can be hidden without changing the optimal action, and what is the canonical exact abstraction that records all and only the distinctions that matter for that choice? We study this question for a decision problem $\mathcal{D}=(A,S,U)$ with $S=X_{1}\times\cdots\times X_{n}$. A coordinate set $I$ is sufficient when agreement on $I$ forces agreement on the optimal-action set:
$s_{I}=s^{\prime}_{I}\implies\text{Opt}(s)=\text{Opt}(s^{\prime}).$
@@ 1 Introduction
The associated optimizer quotient identifies states exactly when they induce the same optimal-action set. In Set, it is the coarsest abstraction through which Opt factors, so every surjective abstraction that preserves optimal-action distinctions factors uniquely through it. This canonical status is theorem-level rather than heuristic. (L: QT1, QT7.)
This paper’s main claim is that once that canonical object is fixed, exact relevance certification becomes a regime-sensitive complexity problem. The structural target is the same throughout, but the cost of certifying it changes with the ambient model. In the static regime, certification reduces to counterexample exclusion and collapses to relevance containment. In the stochastic regime, conditioning splits the exact question into preservation and decisiveness. In the sequential regime, temporal contingency lifts the same certification task to planning-style reasoning. The paper therefore has a simple backbone: the optimizer quotient supplies the canonical notion of exact decision-preserving compression, and the regime analyses classify when that structure can be certified exactly.
Informal interpretive guide.
The optimizer quotient isolates the decision signal: it collapses exactly those state distinctions that do not change the optimal-action set. In that precise sense, it removes exactly decision noise and nothing decision-relevant. Relevant coordinates are points of leverage: changing them can change the decision. In the static regime, the resulting simplification is forced by structure rather than discovered by search. In the later consequence sections, the simplicity tax names the residual burden created when a system uses heuristic simplification in place of exact certification. (L: DN1-2.)
This is a certification problem rather than a forward-evaluation problem. Evaluating $U(a,s)$ or $\text{Opt}(s)$ on a fully specified state is one task; proving that a family of coordinates is irrelevant requires ruling out every counterexample pair of states. That universal structure is exactly what makes the optimizer quotient central rather than cosmetic: it records the maximal lossless collapse of the state space, and every exact decision-preserving abstraction must refine it. In the static regime, this structural picture sharpens further. A coordinate set is sufficient exactly when it exposes all relevant coordinates, so the minimum sufficient set is not discovered by a separate alternating search. It is the relevant-coordinate set itself. By contrast, Anchor-Sufficiency retains a genuine existential choice because the anchor assignment must still satisfy a universal fiberwise condition.
Probability splits the certification problem.
In stochastic decision problems, exact analysis separates into a preservation question and a decisiveness question. Preservation asks whether coarse information reproduces the full-information optimizer, so it is the direct stochastic analogue of exact quotient preservation. Decisiveness asks whether each observable fiber already determines a unique Bayes-optimal action, so it is a stronger decision-completeness demand. Formally, these are stochastic sufficiency and stochastic decisiveness. They are distinct, and the paper treats them separately on purpose. Preservation gives the bridge back to static sufficiency and, under full support, back to the same quotient structure as the static regime, while beyond full support the paper isolates support-sensitive bridge and obstruction lemmas rather than claiming a full general-distribution classification. Decisiveness yields the paper’s strongest succinct-encoding hardness classification. The stochastic row is therefore not one theorem about one predicate, but a split exact theory with two different complexity behaviors.
This is also what the complexity classification buys beyond nearby literatures on reducts, feature selection, or abstraction heuristics. Those traditions often ask whether some smaller representation is informative enough, predictive enough, or value-preserving enough. The present paper asks a sharper question: when can one certify exactly that hiding coordinates leaves the optimal-action correspondence unchanged? That exactness requirement is what later yields a no-general exact minimizer theorem, a regime-sensitive tractability boundary, and a direct reliability corollary about exact certifiers in the hard regime.
Connection to backdoor tractability in CSPs.
Our coordinate sufficiency framework is structurally parallel to the backdoor framework for constraint satisfaction problems [2]. In their setting, a backdoor is a minimal variable set whose fixation collapses a CSP to a tractable class (e.g., one admitting a majority polymorphism). In our setting, a sufficient set is a minimal coordinate set whose observation collapses a decision problem to a determined optimal-action set. Their result that Partition-Majority-CSP is W[2]-hard when the tractable subset is unknown, but fixed-parameter tractable (FPT) when that subset is given, mirrors our result that identifying decision-relevant coordinates is coNP-complete and $\Sigma_{2}^{P}$-complete, while exploiting a known sufficient set is polynomial-time. Both frameworks formalize the same insight: discovering what matters is harder than exploiting it once known.
This distinction also explains why exactness matters. When a system replaces exact relevance certification with a heuristic simplification, the omitted relevance does not disappear; it is shifted into residual uncertainty that must be handled elsewhere in the system. We refer to that externalized burden as the simplicity tax. The results of this paper give that term formal content: exact simplification can be computationally expensive, but inexact simplification does not erase decision-relevant structure; it relocates it.
This systematic development reveals a regime-sensitive complexity landscape because the same structural target is being certified under richer models. Static certification is governed by counterexample exclusion. In the stochastic regime, preservation retains the quotient-preservation target while decisiveness introduces conditional-comparison hardness. In the sequential regime, omitted information can matter only through future contingencies, so exact certification becomes planning-hard. Read this way, the complexity matrix is not a list of unrelated classifications. It is the cost profile of certifying the exact decision-relevant distinctions encoded by the optimizer quotient.
Throughout, we work with finite coordinate-structured decision problems under the encoding conventions fixed in Section 2.4.
1.1 Contributions
The paper is best read as a theory of exact decision-relevant information organized around one canonical mathematical object: the optimizer quotient. The quotient gives the coarsest exact abstraction that preserves optimal-action distinctions, and this canonical status is theorem-level in the artifact. The paper then classifies how the cost of certifying that structure changes across static, stochastic, and sequential regimes, identifies tractable structural subcases, and supports the finite combinatorial core with a Lean artifact.
  1. 1.
    Optimizer quotient as canonical structure. For a finite decision problem, the optimizer quotient identifies states exactly when they induce the same optimal-action set. In Set, it is the coarsest exact decision-preserving abstraction, equivalently the coimage/image factorization of the optimizer map. This gives the paper a common structural object through which the later static, stochastic, and sequential results can be compared.
  2. 2.
    Regime-sensitive certification landscape. The paper studies how hard it is to certify exact preservation of decision-relevant structure across three regimes. In the static regime, sufficiency collapses to relevance containment, so Minimum-Sufficient-Set is coNP-complete rather than the expected $\Sigma_{2}^{P}$ search problem, while Anchor-Sufficiency remains $\Sigma_{2}^{P}$-complete. This collapse is structurally nontrivial. In classical rough-set theory, general decision tables can admit multiple incomparable minimal reducts, which drives the minimum-reduct problem into combinatorial search territory. Under the exact optimal-action preservation predicate studied here, Proposition 2.9 proves there is exactly one minimal sufficient set, the relevant-coordinate set itself. This uniqueness is the structural reason for the collapse to coNP relevance containment. To our knowledge, this specific structural collapse, from $\Sigma_{2}^{P}$ search to coNP driven by optimizer-quotient uniqueness, is a novel classification not captured by general rough-set bounds. In the stochastic regime, exact analysis splits into preservation and decisiveness. Preservation is polynomial-time under explicit-state encoding, bridges back to static sufficiency and the optimizer quotient, and under full support yields inherited coNP and $\Sigma_{2}^{P}$ classifications for the minimum and anchor variants. Decisiveness is polynomial-time in explicit state and PP-hard under succinct encoding, with the anchor and minimum variants PP-hard and in $\textsf{NP}^{\textsf{PP}}$ at paper level. In the sequential regime, sufficiency, minimum, and anchor queries are PSPACE-complete.
  3. 3.
    Tractability islands and encoding-sensitive contrast. The paper proves an explicit-state versus succinct-encoding contrast and isolates tractable subcases across two tiers: simple edge cases (single action, bounded state space, strict dominance, constant optimal set) that reduce to immediate checks, and structurally nontrivial cases (bounded actions, separable utility, low tensor rank, tree structure, bounded treewidth, coordinate symmetry) each removing a distinct source of hardness. These are not ad hoc exceptions. They are regimes in which the decision boundary admits enough structural alignment to make exact certification feasible.
  4. 4.
    Consequences for exact certifiers. In the hard exact regime, certifiers cannot in general be simultaneously sound, complete, and polynomial-budgeted. The same theory yields witness-checking lower bounds, approximation obstructions, and the practical consequence called the simplicity tax: when exact relevance is not certified, the unresolved burden is not removed but shifted elsewhere in the system.
  5. 5.
@@ 1.1 Contributions
    Mechanized finite-decision core. A Lean 4 artifact mechanically checks the optimizer-quotient universal-property core, the main finite deciders, search procedures, reduction-correctness lemmas, witness/checking schemas, and bridge results (L: QT1, QT7, DC91-99, OU1-2, OU8-12, EH1-10.) The oracle-class placements are proved in the paper text, while the artifact independently verifies the finite combinatorial core on which those arguments rely. This places the mechanization in the emerging category of problem-specific certified reduction infrastructure.
1.2 Quantifier Structure
The complexity gap comes from the structure of the question. Insufficiency has short witnesses: two states that agree on the proposed coordinates but induce different optimal-action sets. Sufficiency, by contrast, asks for the absence of every such witness. So the object-level task is forward evaluation on a fully specified state, while the meta-level task is universal verification over counterexample pairs. This difference drives the coNP classification and explains why relevance identification is not just another forward-evaluation problem.
The static regime also contains the paper’s main structural simplification. Although minimum sufficiency can be written with an outer existential quantifier over coordinate sets, the existential layer collapses once sufficiency is characterized as containment of all relevant coordinates. This $\Sigma_{2}^{P}$-to-coNP collapse mirrors a related phenomenon in backdoor tractability for CSPs, where identifying a minimal backdoor to a tractable class is W[2]-hard [2], while exploiting a given backdoor is fixed-parameter tractable. The stochastic and sequential regimes do not simplify in the same way: the stochastic regime separates into preservation and decisiveness, with full-support inheritance on the preservation side and counting-style hardness on the decisiveness side, while temporal evolution introduces dependence on temporally induced optimizer structure.
1.3 Paper Structure
Section 2 introduces the formal setup and encoding conventions. Sections 3, 4, 5, and 6 develop the core theorem package. Sections 7 and 8 give the encoding-sensitive contrast and tractable subcases. Section 9 collects implications for exact simplification and reliability consequences for exact certifiers. Section 10 situates the work relative to formalized complexity and decision theory.
1.4 Artifact Availability
@@
The proof artifact is archived at https://doi.org/10.5281/zenodo.18998870. The cited-content snapshot contains 22068 lines across 67 files with 0 occurrences of sorry.
2 Formal Setup
@@
MINIMUM-SUFFICIENT-SET apparent $\exists I\,\forall(s,s^{\prime})$ check smallest sufficient support collapses to relevance containment in the static regime
@@
Setup map. The four central objects differ mainly in quantifier shape: minimum sufficiency collapses in the static regime, while anchor sufficiency keeps its existential fiber choice.
2.1 Decision Problems with Coordinate Structure
Definition 2.1 (Decision Problem).
@@ 2.1 Decision Problems with Coordinate Structure
is the projection of $s$ onto coordinates in $I$.
Definition 2.3 (Optimizer Map).
@@
$\text{Opt}(s):=\arg\max_{a\in A}U(a,s)=\{a\in A:U(a,s)=\max_{a^{\prime}\in A}U(a^{\prime},s)\}$
2.2 Sufficiency and Relevance
Definition 2.4 (Sufficient Coordinate Set).
@@ 2.2 Sufficiency and Relevance
$\forall s,s^{\prime}\in S:\quad s_{I}=s^{\prime}_{I}\implies\text{Opt}(s)=\text{Opt}(s^{\prime})$
Equivalently, the optimal action depends only on coordinates in $I$.
Proposition 2.5 (Empty-Set Sufficiency Equals Constant Decision Boundary).
The empty set $\emptyset$ is sufficient if and only if $\text{Opt}(s)$ is constant over the entire state space.
(L: DP6.)
@@
If $\emptyset$ is sufficient, then every two states agree on their empty projection, so sufficiency forces $\text{Opt}(s)=\text{Opt}(s^{\prime})$ for all $s,s^{\prime}\in S$. The converse is immediate.  $\blacksquare$
Proposition 2.6 (Insufficiency Equals Counterexample Witness).
Coordinate set $I$ is not sufficient if and only if there exist states $s,s^{\prime}\in S$ such that $s_{I}=s^{\prime}_{I}$ but $\text{Opt}(s)\neq\text{Opt}(s^{\prime})$.
(L: DP7.)
@@
This is the direct negation of Definition 2.4.  $\blacksquare$
Definition 2.7 (Minimal Sufficient Set).
A sufficient set $I$ is minimal if no proper subset $I^{\prime}\subsetneq I$ is sufficient.
Definition 2.8 (Relevant Coordinate).
@@
$i\text{ is relevant}\iff\exists s,s^{\prime}\in S:\;\Big(\forall j\neq i,\;s_{j}=s^{\prime}_{j}\Big)\ \wedge\ \text{Opt}(s)\neq\text{Opt}(s^{\prime}).$
Proposition 2.9 (Minimal-Set/Relevance Equivalence).
For any minimal sufficient set $I$ and any coordinate $i$,
@@
If $i\in I$ were not relevant, then changing coordinate $i$ while keeping all others fixed would never alter the optimal-action set, so removing $i$ would preserve sufficiency, contradicting minimality. Conversely, every sufficient set must contain each relevant coordinate, because omitting a relevant coordinate leaves a witness pair with identical $I$-projection but different optimal-action sets.  $\blacksquare$
Definition 2.10 (Exact Relevance Identifiability).
For a decision problem $\mathcal{D}$ and candidate set $I$, we say that $I$ is exactly relevance-identifying if
@@
Definition 2.11 (Structural Rank).
@@
2.3 Optimizer Quotient
Definition 2.12 (Decision Equivalence).
@@ 2.3 Optimizer Quotient
$s\sim_{\text{Opt}}s^{\prime}\iff\text{Opt}(s)=\text{Opt}(s^{\prime}).$
Definition 2.13 (Decision Quotient).
@@
Proposition 2.14 (Optimizer Quotient as Coimage/Image Factorization).
The decision quotient $Q_{\mathcal{D}}$ is canonically in bijection with $\operatorname{range}(\text{Opt})$. Equivalently, in Set, it is the coimage/image factorization of the optimizer map.
(L: DQ6-7.)
@@
Two states are identified by $\sim_{\text{Opt}}$ exactly when they have the same optimizer value. Hence quotient classes are in one-to-one correspondence with attained values of Opt.  $\blacksquare$
Theorem 2.15 (Universal Property of the Optimizer Quotient).
Let $Q_{\mathcal{D}}=S/{\sim_{\text{Opt}}}$ be the decision quotient with quotient map $\pi:S\to Q_{\mathcal{D}}$. For any surjective abstraction $\phi:S\to T$ such that
$\phi(s)=\phi(s^{\prime})\implies\text{Opt}(s)=\text{Opt}(s^{\prime}),$
@@
Because $\phi$ preserves optimizer values on its fibers, any two states identified by $\phi$ are decision-equivalent. So each fiber of $\phi$ is contained in a single $\sim_{\text{Opt}}$-class.
For each $t\in T$, choose any $s\in S$ with $\phi(s)=t$ and define $\psi(t):=\pi(s)$. This is well defined because any two such choices lie in the same $\sim_{\text{Opt}}$-class. By construction, $(\psi\circ\phi)(s)=\pi(s)$ for every $s\in S$, so $\pi=\psi\circ\phi$. Uniqueness follows from surjectivity of $\phi$: if $\psi^{\prime}$ also satisfies $\pi=\psi^{\prime}\circ\phi$, then every $t\in T$ has the form $t=\phi(s)$ and therefore
@@
Remark 2.16 (Canonical Status).
@@
(L: QT7, AB2.)
Remark 2.17 (Decision Noise).
@@
(L: DN1-2, DN5.)
Proposition 2.18 (Sufficiency Characterization).
Coordinate set $I$ is sufficient if and only if the optimizer map factors through the projection $\pi_{I}:S\to S_{I}$. Equivalently, there exists a function $\overline{\text{Opt}}_{I}:S_{I}\to\mathcal{P}(A)$ such that
$\text{Opt}=\overline{\text{Opt}}_{I}\circ\pi_{I}.$
(L: DQ42.)
@@
If $I$ is sufficient, define $\overline{\text{Opt}}_{I}(x)$ to be $\text{Opt}(s)$ for any state $s$ with $s_{I}=x$; this is well defined exactly because sufficiency makes Opt constant on each projection fiber. Conversely, any such factorization forces states with equal $I$-projection to have equal optimal-action sets.  $\blacksquare$
2.4 Computational Model and Input Encoding

### 31 Mar 2026 v2

# The Optimizer Quotient and the Certification Trilemma

@@ Abstract
The optimizer quotient is the canonical object for exact decision-relevant information: it is the coarsest exact decision-preserving abstraction (Theorem 2.15). This paper proves that exact certification of this object’s coordinate structure is subject to an impossibility trilemma: under $\mathrm{P}\neq\mathrm{coNP}$, no certifier can be simultaneously sound, complete on all in-scope instances, and polynomial-budgeted (Theorem 7.1). The cost of this impossibility varies by regime: coNP (static), $P$-hard (stochastic decisiveness), PSPACE-complete (sequential). Six structural restrictions collapse certification to polynomial time. The finite reduction and verification core is mechanized in Lean 4.
I Introduction
I-A The Optimizer Quotient: A Canonical Object
@@
Exact decision-preserving information has a canonical form: the optimizer quotient. For a finite coordinate-structured decision problem $\mathcal{D}=(A,S,U)$ with $S=X_{1}\times\cdots\times X_{n}$, the optimizer quotient identifies states exactly when they induce the same optimal-action set $\mathrm{Opt}(s)$. This quotient is the coarsest exact decision-preserving abstraction (Theorem II.15): every surjective abstraction that preserves optimal actions factors uniquely through it. Equivalently, in Set, it is the coimage/image factorization of the optimizer map.
This canonical object exists for every decision problem and unifies separate literatures. In rough-set theory, attribute reduction asks which coordinates preserve decision distinctions; in constraint satisfaction, backdoor sets identify variables whose fixation yields tractability; in feature selection, minimal sufficient subsets are sought; in MDP abstraction, state aggregation preserves value or policy structure. The optimizer quotient provides a common mathematical foundation: it is the canonical object for exact relevance in all these settings.
I-B The Certification Trilemma: An Impossibility Theorem
From the optimizer quotient we derive a structural constraint on exact certifiers. Under $\mathrm{P}\neq\mathrm{coNP}$, no solver can be simultaneously integrity-preserving, non-abstaining on all in-scope instances, and polynomial-budgeted (Theorem 7.1). This impossibility yields a trilemma: any exact certifier in the hard regime must either abstain on some instances, weaken its guarantee, or risk overclaiming. The trilemma is not merely a restatement of coNP-hardness; it is a certified-solver impossibility result with operational consequences for systems claiming exact reliability.
I-C The Complexity Map as Evidence
The impossibility manifests differently across regimes. The static regime yields coNP-completeness for base and minimum queries (via relevance containment), while static anchor sufficiency is $\Sigma_{2}^{P}$-complete. In the stochastic regime, the problem splits: preservation bridges to static sufficiency (polynomial-time under explicit encoding), while decisiveness is $P$-hard under succinct encoding. In the sequential regime, temporal contingency lifts all queries to PSPACE-complete. These classifications form a regime-sensitive complexity matrix (Table III) that shows the impossibility’s reach expanding with model expressiveness. Two narrow cells in the stochastic-preservation row (general distributions, without the full-support assumption) remain open; these are explicitly marked as scope boundaries in Table II and do not affect any of the proved results, the optimizer quotient, the certification trilemma, or the completeness of the static, sequential, and stochastic-decisiveness regimes.
I-D Tractable Boundaries
Six structural restrictions collapse exact certification to polynomial time: bounded actions, separable utility, low tensor rank, tree structure, bounded treewidth, and coordinate symmetry. Each removes a distinct source of hardness, providing tractable boundaries to the impossibility theorem.
I-E Mechanized Core
@@
A Lean 4 artifact mechanically checks the optimizer-quotient universal property, finite deciders, reduction-correctness lemmas, witness schemas, and bridge results  [^1] Lean: QT1, QT7, DC91-99, OU1-2, OU8-12, EH1-10. The 22K-line artifact verifies the combinatorial core, while oracle-class placements are argued in the paper text.
I-F Paper Structure
Section II introduces the formal setup. Sections III, IV, and V develop the static, stochastic, and sequential complexity classifications. Section VI consolidates them into the regime matrix. Section 7 presents the impossibility theorem and trilemma. Section 8 collects structural consequences. Sections VII and VIII cover encoding contrasts and tractable cases. Section 9 situates the work, and Appendix B collects model translations.
I-G Artifact Availability
@@
The proof artifact is archived at https://doi.org/10.5281/zenodo.19057595. The cited-content snapshot contains 22068 lines across 67 files with 0 occurrences of sorry.
II Formal Setup
@@
MINIMUM-SUFFICIENT-SET apparent $\exists I\,\forall(s,s^{\prime})$ check smallest sufficient support collapses to relevance containment in static regime
@@
TABLE I: Setup map. The four central objects differ mainly in quantifier shape: minimum sufficiency collapses in the static regime, while anchor sufficiency keeps its existential fiber choice.
II-A Decision Problems with Coordinate Structure
Definition II.1 (Decision Problem).
@@ II-A Decision Problems with Coordinate Structure
is the projection of $s$ onto coordinates in II.
Definition II.3 (Optimizer Map).
@@
$\mathrm{Opt}(s):=\arg\max_{a\in A}U(a,s)=\{a\in A:U(a,s)=\max_{a^{\prime}\in A}U(a^{\prime},s)\}$
II-B Sufficiency and Relevance
Definition II.4 (Sufficient Coordinate Set).
@@ II-B Sufficiency and Relevance
$\forall s,s^{\prime}\in S:\quad s_{I}=s^{\prime}_{I}\implies\mathrm{Opt}(s)=\mathrm{Opt}(s^{\prime})$
Equivalently, the optimal action depends only on coordinates in II.
Proposition II.5footnotemark: footnotetext: : DP6 (Empty-Set Sufficiency Equals Constant Decision Boundary).
The empty set $\emptyset$ is sufficient if and only if $\mathrm{Opt}(s)$ is constant over the entire state space.
@@
If $\emptyset$ is sufficient, then every two states agree on their empty projection, so sufficiency forces $\mathrm{Opt}(s)=\mathrm{Opt}(s^{\prime})$ for all $s,s^{\prime}\in S$. The converse is immediate.  $\blacksquare$
Proposition II.6footnotemark: footnotetext: : DP7 (Insufficiency Equals Counterexample Witness).
Coordinate set II is not sufficient if and only if there exist states $s,s^{\prime}\in S$ such that $s_{I}=s^{\prime}_{I}$ but $\mathrm{Opt}(s)\neq\mathrm{Opt}(s^{\prime})$.
@@
This is the direct negation of Definition II.4.  $\blacksquare$
Definition II.7 (Minimal Sufficient Set).
A sufficient set II is minimal if no proper subset $I^{\prime}\subsetneq I$ is sufficient.
Definition II.8 (Relevant Coordinate).
@@
$i\text{ is relevant}\iff\exists s,s^{\prime}\in S:\;\Big(\forall j\neq i,\;s_{j}=s^{\prime}_{j}\Big)\ \wedge\ \mathrm{Opt}(s)\neq\mathrm{Opt}(s^{\prime}).$
Proposition II.9footnotemark: footnotetext: : DP1-2 (Minimal-Set/Relevance Equivalence).
For any minimal sufficient set II and any coordinate $i$,
@@
If $i\in I$ were not relevant, then changing coordinate $i$ while keeping all others fixed would never alter the optimal-action set, so removing $i$ would preserve sufficiency, contradicting minimality. Conversely, every sufficient set must contain each relevant coordinate, because omitting a relevant coordinate leaves a witness pair with identical II-projection but different optimal-action sets.  $\blacksquare$
Definition II.10 (Exact Relevance Identifiability).
For a decision problem $\mathcal{D}$ and candidate set II, we say that II is exactly relevance-identifying if
@@
Definition II.11 (Structural Rank).
@@
II-C Optimizer Quotient
Definition II.12 (Decision Equivalence).
@@ II-C Optimizer Quotient
$s\sim_{\mathrm{Opt}}s^{\prime}\iff\mathrm{Opt}(s)=\mathrm{Opt}(s^{\prime}).$
Definition II.13 (Decision Quotient).
@@
Proposition II.14footnotemark: footnotetext: : DQ6-7 (Optimizer Quotient as Coimage/Image Factorization).
The decision quotient $Q_{\mathcal{D}}$ is canonically in bijection with $\operatorname{range}(\mathrm{Opt})$. Equivalently, in Set, it is the coimage/image factorization of the optimizer map.
@@
Two states are identified by $\sim_{\mathrm{Opt}}$ exactly when they have the same optimizer value. Hence quotient classes are in one-to-one correspondence with attained values of $\mathrm{Opt}$.  $\blacksquare$
Theorem II.15footnotemark: footnotetext: : QT1, QT7 (Universal Property of the Optimizer Quotient).
Let $Q_{\mathcal{D}}=S/{\sim_{\mathrm{Opt}}}$ be the decision quotient with quotient map $\pi:S\to Q_{\mathcal{D}}$. For any surjective abstraction $\phi:S\to T$ such that
$\phi(s)=\phi(s^{\prime})\implies\mathrm{Opt}(s)=\mathrm{Opt}(s^{\prime}),$
@@
Because $\phi$ preserves optimizer values on its fibers, any two states identified by $\phi$ are decision-equivalent. So each fiber of $\phi$ is contained in a single $\sim_{\mathrm{Opt}}$-class.
For each $t\in T$, choose any $s\in S$ with $\phi(s)=t$ and define $\psi(t):=\pi(s)$. This is well defined because any two such choices lie in the same $\sim_{\mathrm{Opt}}$-class. By construction, $(\psi\circ\phi)(s)=\pi(s)$ for every $s\in S$, so $\pi=\psi\circ\phi$. Uniqueness follows from surjectivity of $\phi$: if $\psi^{\prime}$ also satisfies $\pi=\psi^{\prime}\circ\phi$, then every $t\in T$ has the form $t=\phi(s)$ and therefore
@@
Remark II.16footnotemark: footnotetext: : QT7, AB2. (Canonical Status).
@@
Remark II.17footnotemark: footnotetext: : DN1-2, DN5. (Decision Noise).
@@
Proposition II.18footnotemark: footnotetext: : DQ42 (Sufficiency Characterization).
Coordinate set II is sufficient if and only if the optimizer map factors through the projection $\pi_{I}:S\to S_{I}$. Equivalently, there exists a function $\overline{\mathrm{Opt}}_{I}:S_{I}\to\mathcal{P}(A)$ such that
$\mathrm{Opt}=\overline{\mathrm{Opt}}_{I}\circ\pi_{I}.$
@@
If II is sufficient, define $\overline{\mathrm{Opt}}_{I}(x)$ to be $\mathrm{Opt}(s)$ for any state $s$ with $s_{I}=x$; this is well defined exactly because sufficiency makes $\mathrm{Opt}$ constant on each projection fiber. Conversely, any such factorization forces states with equal II-projection to have equal optimal-action sets.  $\blacksquare$
II-D Computational Model and Input Encoding